Jason and Hilger are required to paint over the graffiti on a wall.
If Jason worked alone, it would take him 20 hrs to repaint. Working
alone, Hilger could do the job in 15 hrs. How long will it take them
to do the painting If they work together?

Found 2 solutions by Edwin McCravy, ikleyn:
Answer by Edwin McCravy(20054)

Jason and Hilger are required to paint over the graffiti on a wall.
If Jason worked alone, it would take him 20 hrs to repaint. Working
alone, Hilger could do the job in 15 hrs. How long will it take them
to do the painting If they work together?

Jason's working rate is 1 job in 20 hrs or 1/20 jobs per hour.
Hilger's working rate is 1 job in 15 hrs or 1/15 jobs per hour.

Their combined rate = 1/20 + 1/15 = 3/60 + 4/60 = 7/60 jobs per hour.

(Number of jobs done) = (rate in number of jobs per hour) × (number of hours)

Number of jobs done = 1 job
Rate in number of jobs per hour = 7/60
Number of hours = x

(Number of jobs done) = (rate in number of jobs per hour) × (number of hours

                    1 = (7/60)(x)

Multiply both sides by 60 and solve for x

Edwin

Answer by ikleyn(52778)

.
This problem is a typical joint work problem.

For a wide variety of similar solved joint-work problems with detailed explanations see the lessons
    - Using Fractions to solve word problems on joint work,
    - Solving more complicated word problems on joint work,
    - Selected joint-work word problems from the archive
in this site.

Read them and get be trained in solving joint-work problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the topic "Rate of work and joint work problems" of the section "Word problems".